OpenStudy (anonymous):

please help! The formula for summing an arithmetic progression is S= 0.5n(2a+(n-1)d) if a = 8, n = 20 and S = 225, find d. please explain step by step the answer is d=0.5

OpenStudy (dayakar):

s= 0.5n [2a+(n-1)d] 225 = 0.5*20[2*8+(20-1) d] can u find d now?

OpenStudy (dayakar):

do u understand it!!!!!!!!!!!?

OpenStudy (anonymous):

just use pemdas

OpenStudy (anonymous):

@maryanb

OpenStudy (jjcoolja12):

Solution 1 Simplify 20−120−1 to 1919 225=0.5×20(2×8+19d) 225=0.5×20(2×8+19d) 2 Simplify 2×82×8 to 1616 225=0.5×20(16+19d) 225=0.5×20(16+19d) 3 Simplify 0.5×20(16+19d)0.5×20(16+19d) to 10(16+19d)10(16+19d) 225=10(16+19d) 225=10(16+19d) 4 Divide both sides by 1010 22510=16+19d 22510=16+19d 5 Simplify 2251022510 to 452452 452=16+19d 452=16+19d 6 Subtract 1616 from both sides 452−16=19d 452−16=19d 7 Simplify 452−16452−16 to 132132 132=19d 132=19d 8 Divide both sides by 1919 132×19=d 132×19=d 9 Simplify 2×192×19 to 3838 1338=d 1338=d 10 Switch sides d=13/38

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got this or no? explanation above must be from some web site